OpenStudy (anonymous):

Show all work so I can see how it's done please. Find the following product: (y-5)^2

OpenStudy (anonymous):

Use the FOIL method.

OpenStudy (amistre64):

1+2+1 and fill in the as and bs ... makes sense in me head

OpenStudy (anonymous):

(y-5)^2 = (y-5)(y-5) = y^2 -5y - 5y + 25 = y^2 - 10y + 25

OpenStudy (anonymous):

\[(y-5)^2\] \[y^2 - 5y - 5y + 25\] \[y^2 - 10y +25\]

OpenStudy (amistre64):

FOIL is just memory clutter; you already know how to multiply. y-5 y-5 ---- y^2 -5y -5y +25 ------------- y^2 -10y + 25

OpenStudy (anonymous):

multiply the terms in second bracket by y then do the same with -5

OpenStudy (anonymous):

i agree amistre - there is no real need for FOIL

OpenStudy (anonymous):

amistre, how did you write your solution?

OpenStudy (amistre64):

y-5 times y-5 ----- y.y-5y -5y+25 ----------- or the more normal way of multiplying may be: y-5 times y-5 ----- -5y+25 y.y-5y -----------

OpenStudy (amistre64):

with polys i like to start at the left and go right

OpenStudy (nubeer):

(y-5)^2 (y)^2-2(y)(5)+(5)^2 y^2-10y+25

OpenStudy (anonymous):

whoops nvm, i thought for a second you were using a program to write/draw your solution. but your typing :P my bad.

OpenStudy (amistre64):

7*5 (3+4) (2+3) 3 + 4 2 + 3 ------ 3.2+4.2 +3.3 +4.3 ------------ 6 + 8 + 9 + 12 = 35

OpenStudy (amistre64):

lol ... good old ascii characters ;)